# A Survey of Reinforcement Learning For Economics

### 17 Mar 2026 v3

@@ 4.1.6 SARSA (1994)
If the agent follows an $\varepsilon$-greedy policy, SARSA converges to $Q^{\varepsilon\text{-greedy}}$, not $Q^{*}$ [^32] SARSA converges to $Q^{*}$ under GLIE (greedy in the limit with infinite exploration)A GLIE schedule explores all state-action pairs infinitely often but converges to the greedy policy asymptotically. The standard $\varepsilon$-greedy policy with $\varepsilon_{t}\to 0$ is one such schedule. policies and standard step-size conditions (Section 5.2). This distinction matters when exploration is costly. Consider the cliff-walking problem, where an agent must traverse a gridworld with a cliff along one edge. The optimal path runs along the cliff edge (shortest route), but the $\varepsilon$-greedy policy occasionally falls off. Q-learning learns the optimal path because it evaluates the greedy policy; the agent falls off during learning but the Q-values reflect the optimal route. SARSA learns a safer path further from the cliff because it evaluates the actual exploratory policy; it accounts for the fact that exploration sometimes leads to catastrophic states.
@@ 5.2.8 Rollout, Lookahead, and AlphaZero
Recall the AlphaGo Zero system from Section 4.2.4, where a neural network $f_{\theta}(s)=(\mathbf{p},v)$ outputs a prior policy $\mathbf{p}$ and a value estimate $v$ for any board position $s$. During play, the network does not act alone: Monte Carlo Tree Search runs simulated games from the current position, using $v$ to evaluate leaf nodes and $\mathbf{p}$ to guide which branches to explore. [^73] AlphaGo Zero uses 1,600 MCTS simulations per move for Go; the generalized AlphaZero algorithm uses 800 simulations per move across chess, shogi, and Go (Silver et al., 2018, Table S3). The network provides $\tilde{V}$; MCTS applies multi-step lookahead through selective tree expansion. Table 7 makes the correspondence explicit.
@@ 5.4.3 Natural Policy Gradient and Gradient Domination
The RL objective $J(\theta)$ is non-convex in $\theta$. For economists trained to distrust gradient methods on non-convex objectives, the natural concern is convergence to spurious local optima. For tabular softmax policies (one free parameter per state-action pair), this concern is unfounded. The landscape is “benign” in a precise sense. Agarwal et al. (2021a) prove that $J(\theta)$ satisfies a gradient domination condition (also called Polyak-Łojasiewicz, or PL). The PL condition has the same functional form as the strong convexity condition for guaranteeing linear convergence of gradient descent, but it applies to non-convex functions: whenever $\|\nabla J(\theta)\|$ is small, the policy must be near-optimal. Formally, the sub-optimality $J(\pi^{*})-J(\pi_{\theta})$ is bounded by a constant times $\|\nabla J(\theta)\|^{2}$. The implication is immediate: any point where the gradient vanishes is globally optimal. The non-convex landscape has no false peaks, no spurious local maxima. Gradient ascent cannot get trapped.
@@ 5.5.1 Actor-Critic Architecture and Two-Timescale Convergence
Convergence requires a structural condition on the critic. The critic’s feature vectors must span the actor’s score functions $\nabla_{\omega}\log\pi_{\omega}(a|s)$, so that the critic’s approximation error lies orthogonal to the policy gradient direction. Under this compatibility condition, the critic’s projection error does not bias the actor’s gradient estimates. [^92] The compatible function approximation theorem first appears in Sutton et al. (2000) and is the key structural requirement in Konda and Tsitsiklis (2000). It constrains the critic architecture to be “compatible” with the actor parameterization, the same condition that makes the natural policy gradient equal to the critic’s weight vector in Kakade (2002).
A2C (Advantage Actor-Critic) is the synchronous variant: collect a batch of transitions, compute TD errors, and update both networks. A3C (Mnih et al., 2016) parallelizes this across multiple workers updating a shared parameter server asynchronously. [^93] Parallel workers decorrelate gradient estimates by exploring different parts of the state space simultaneously, removing the need for an experience replay buffer. Rigorous convergence theory for A3C’s lock-free asynchronous parameter updates remains an open problem; existing analyses of asynchronous stochastic approximation (Qu and Wierman, 2020) address classical asynchrony (different state-action pairs updated at different times on a single trajectory), not the parallel-worker gradient setting.
@@ 5.5.2 Entropy Regularization and Soft Actor-Critic
where $\mathcal{H}(\pi)=-\sum_{a}\pi(a)\log\pi(a)$ is the entropy and $\tau>0$ is the temperature parameter. Geist et al. (2019) later provided the unifying theoretical framework, showing that entropy regularization converts the Bellman optimality operator’s non-smooth hard max into a smooth log-sum-exp, and that the resulting soft Bellman operator remains a $\gamma$-contraction, preserving the convergence guarantees of standard dynamic programming. [^94] The optimal policy under entropy regularization is $\pi^{*}(a|s)\propto\exp(Q^{*}(s,a)/\tau)$, which is precisely the McFadden (1974) logit choice probability with systematic utility $Q^{*}(s,a)$ and scale parameter $\tau$. The entropy-regularized value function is the log-sum-exp of Q-values, corresponding to the inclusive value (log-sum) operator in nested logit models. This equivalence between the soft-control framework and dynamic discrete choice models with EV1 taste shocks is developed formally in Rust and Rawat (2026), Appendix A.
@@ 5.5.3 Error Amplification Under Approximate Value Functions
At $\gamma=0.99$, the amplification factor is $2\cdot 0.99/0.01=198$. A 1% error in value function approximation yields at most 198% error in policy value. [^95] The Singh-Yee bound is worst-case and not tight in general; massoud Farahmand et al. (2010) derive tighter bounds under smoothness assumptions on the MDP. The amplification factor $2\gamma/(1-\gamma)$ is a sensitivity analysis: it quantifies how errors in the “inputs” (value estimates) propagate to “outputs” (policy quality), analogous to errors-in-variables bias in regression. The discount factor $\gamma$ controls sensitivity; more patient agents face larger amplification. This bound is pessimistic but finite: approximate value functions do not cause unbounded policy degradation.
@@ 5.5.4 Sample Complexity of Planning
Classical dynamic programming complexity scales with the state space size $|\mathcal{S}|$. For problems like Go, where $|\mathcal{S}|\approx 10^{170}$, exact computation is impossible. Kearns et al. (2002) prove that with access to a generative model [^96] A “generative model” in RL is a simulator that, given any state-action pair $(s,a)$, returns a sampled next state $s^{\prime}\sim P(\cdot|s,a)$ and reward $r$. This is unrelated to “generative models” in machine learning (GANs, diffusion models) or “generative processes” in Bayesian econometrics. The distinction matters: planning with a generative model is strictly easier than learning from a single trajectory, because the agent can query arbitrary states rather than following a sequential path. (a simulator that samples transitions from any state-action pair), near-optimal planning is possible with no dependence on $|\mathcal{S}|$. The cost is exponential dependence on the effective horizon $H=\log(R_{\max}/(\epsilon(1-\gamma)))/\log(1/\gamma)$: the sparse sampling algorithm requires $O((|\mathcal{A}|/\epsilon)^{H})$ simulator calls. [^97] The sparse sampling algorithm builds a random tree of depth $H$ from the current state, sampling $C$ successor states per action at each node, then estimates values by averaging leaf rewards back up the tree. Its running time is $O((C|\mathcal{A}|)^{H})$, which is exponential in $H$ but entirely independent of $|\mathcal{S}|$. Kearns et al. (2002) also prove a lower bound of $\Omega(2^{H})$ generative model calls for any planning algorithm, so exponential horizon dependence is unavoidable in the worst case. For $\gamma$ near 1, $H\approx(1-\gamma)^{-1}\log(R_{\max}/\epsilon)$, so the method is practical only for short effective horizons or moderate discount factors. The key insight is the tradeoff: classical DP scales linearly in $|\mathcal{S}|$ but polynomially in $H$; sparse sampling eliminates state-space dependence at the cost of exponential horizon dependence. This explains why MCTS succeeds in large state spaces with bounded lookahead.
@@
The minimax-optimal sample complexity for planning with a generative model, when queries to arbitrary state-action pairs are permitted, is $\Theta(|\mathcal{S}||\mathcal{A}|/((1-\gamma)^{3}\epsilon^{2}))$ (Azar et al., 2013). This bound scales linearly in $|\mathcal{S}|$ but polynomially in $1/(1-\gamma)$, the opposite regime from sparse sampling. Agarwal et al. (2020a) show that the plug-in model-based approach (learn $\hat{P}$ from samples, then plan with $\hat{P}$) achieves this minimax rate, establishing that model-based RL is statistically optimal. Li et al. (2024b) further tighten this result by breaking the $|\mathcal{S}||\mathcal{A}|/(1-\gamma)^{2}$ sample-size barrier, showing that the minimax rate is achievable with total sample size as low as $|\mathcal{S}||\mathcal{A}|/(1-\gamma)$. [^98] Recent extensions push these results beyond standard MDPs. Clavier et al. (2024) study the robust MDP setting where the agent must plan under model uncertainty with sa-rectangular or s-rectangular uncertainty sets, establishing minimax rates for robust policy optimization. Wang et al. (2025) extend the analysis to risk-sensitive objectives under the iterated CVaR criterion.

### 23 Mar 2026 v4

@@ 2.3 Structural Equivalences
These equivalences reflect a common convex-analytic structure. The log-sum-exp value function and the negative Shannon entropy are Fenchel conjugates of each other, so maximizing expected payoffs with an entropy bonus and recovering utilities from observed choice probabilities are two views of the same optimization. Chiong et al. (2016) build on this conjugate duality to develop identification and estimation methods for dynamic discrete choice models, and Fosgerau and Sørensen (2022) show that the underlying structure extends well beyond the logit case.
The two fields arrived at this shared mathematics from opposite directions. In RL, entropy regularization began as a pragmatic trick to prevent premature convergence and encourage exploration (Williams and Peng, 1991), and only decades later did Ziebart (2010) and Levine (2018) provide decision-theoretic foundations showing that the resulting policies are robust to model misspecification. In economics, the softmax emerged not from any concern about exploration but from the random utility framework of McFadden (1974), where agents are fully informed and choose deterministically; the apparent randomness arises entirely from the econometrician’s inability to observe all relevant taste variation. The RL agent randomizes because it is ignorant of the environment; the economic agent appears to randomize because the observer is ignorant of the agent’s preferences.
@@ 4.1.6 SARSA (1994)
If the agent follows an $\varepsilon$-greedy policy, SARSA converges to $Q^{\varepsilon\text{-greedy}}$, not $Q^{*}$ [^34] SARSA converges to $Q^{*}$ under GLIE (greedy in the limit with infinite exploration). A GLIE schedule explores all state-action pairs infinitely often but converges to the greedy policy asymptotically. The standard $\varepsilon$-greedy policy with $\varepsilon_{t}\to 0$ is one such schedule. policies and standard step-size conditions (Section 5.2). This distinction matters when exploration is costly. Consider the cliff-walking problem, where an agent must traverse a gridworld with a cliff along one edge. The optimal path runs along the cliff edge (shortest route), but the $\varepsilon$-greedy policy occasionally falls off. Q-learning learns the optimal path because it evaluates the greedy policy; the agent falls off during learning but the Q-values reflect the optimal route. SARSA learns a safer path further from the cliff because it evaluates the actual exploratory policy; it accounts for the fact that exploration sometimes leads to catastrophic states.
@@ 4.2.3 Soft Actor-Critic (2018)
4.2.4 Control as Probabilistic Inference
The control-as-inference framework (Todorov, 2006; Kappen, 2011; Ziebart, 2010; Levine, 2018) recasts the preceding algorithms as instances of probabilistic inference (computation, not statistical inference) in a single graphical model. [^42] The framework is a mathematical language, much like random utility in econometrics, that reveals structural connections rather than deriving algorithms from first principles. The payoff is that every advance in approximate inference (variational methods, message passing, amortized inference) becomes a candidate RL algorithm, and the forward problem (finding the optimal policy given rewards) and the inverse problem (recovering rewards from observed behavior) become two queries in the same model. The construction introduces a binary “optimality” variable $\mathcal{O}_{t}\in\{0,1\}$ at each time step, appended to the standard state-action-dynamics chain (Figure 2). The distribution over this variable is defined as
$P(\mathcal{O}_{t}=1\mid s_{t},a_{t})=\exp(r(s_{t},a_{t})/\tau)$ (29)
where $\tau>0$ is a temperature parameter and rewards satisfy $r\leq 0$. [^43] The non-positive reward assumption ensures $\exp(r/\tau)\in(0,1]$. Any bounded reward can be shifted to satisfy this without changing the optimal policy. Alternatively, the optimality variable can be replaced by an undirected potential $\Phi(s_{t},a_{t})=\exp(r(s_{t},a_{t})/\tau)$, yielding a conditional random field formulation that removes this restriction (Ziebart, 2010). This is a modeling assumption, not a derivation from first principles; the exponential form is chosen so the resulting posterior matches entropy-regularized RL.
[…]
Figure 2: The control-as-inference graphical model. States $s_{t}$ and actions $a_{t}$ jointly determine the next state $s_{t+1}$ (dynamics) and the optimality variable $\mathcal{O}_{t}$ (reward signal). Optimality nodes $\mathcal{O}_{t}$ are observed as $\mathcal{O}_{t}=1$; the posterior over actions $P(a_{t}\mid s_{t},\mathcal{O}_{t:T}=1)$ yields the optimal policy.
The RL problem becomes a probabilistic inference query. Given that all future time steps are optimal ($\mathcal{O}_{t:T}=1$), what is $P(a_{t}\mid s_{t},\mathcal{O}_{t:T}=1)$? Define backward messages $\beta_{t}(s,a)=P(\mathcal{O}_{t:T}=1\mid s_{t}=s,\,a_{t}=a)$, the probability that everything from $t$ onward is optimal given the agent is in state $s$ taking action $a$. These are computed by backward recursion:
$\beta_{T}(s,a)$ $=\exp(r(s,a)/\tau)$ (30)
$\beta_{t}(s,a)$ $=\exp(r(s,a)/\tau)\sum_{s^{\prime}}P(s^{\prime}|s,a)\,\beta_{t+1}(s^{\prime})\quad(t<T)$ (31)
where $\beta_{t+1}(s^{\prime})\propto\sum_{a^{\prime}}\beta_{t+1}(s^{\prime},a^{\prime})$ marginalizes over future actions under a uniform prior. By Bayes’ rule, the posterior over actions is $P(a_{t}\mid s_{t},\mathcal{O}_{t:T}=1)=\beta_{t}(s_{t},a_{t})\,/\,\sum_{a^{\prime}}\beta_{t}(s_{t},a^{\prime})$. Defining $Q(s,a)=\tau\log\beta_{t}(s,a)$, so that backward messages in log-space correspond to soft value functions, and $V(s)=\tau\log\sum_{a}\exp(Q(s,a)/\tau)$ [^44] Since $\beta_{t}(s)\propto\sum_{a}\beta_{t}(s,a)=\sum_{a}\exp(Q(s,a)/\tau)$, we have $V(s)=\tau\log\sum_{a}\exp(Q(s,a)/\tau)$., the posterior becomes $\pi(a\mid s)=\exp\bigl((Q(s,a)-V(s))/\tau\bigr)$, the softmax over Q-values.
Exact inference in this graphical model under stochastic dynamics produces risk-seeking policies. The backup becomes $Q(s,a)=r(s,a)+\tau\log\mathbb{E}_{s^{\prime}}[\exp(V(s^{\prime})/\tau)]$, which overweights unlikely favorable transitions because the agent implicitly assumes it can influence the dynamics. [^45] Under exact inference, the posterior dynamics $P(s^{\prime}|s,a,\mathcal{O}_{1:T}=1)$ differ from the true dynamics $P(s^{\prime}|s,a)$, biasing the inferred transitions toward favorable outcomes. The log-expectation-of-exponentials exceeds the expectation by Jensen’s inequality. This optimism is an artifact of exact inference in the model, not a feature. The framework corrects this by applying structured variational inference, restricting the variational family to distributions that match the true dynamics $P(s^{\prime}|s,a)$. This yields the soft Bellman equations
$Q(s,a)$ $=r(s,a)+\gamma\,\mathbb{E}_{s^{\prime}\sim P}[V(s^{\prime})]$ (32)
$V(s)$ $=\tau\log\sum_{a\in\mathcal{A}}\exp\!\left(\frac{Q(s,a)}{\tau}\right)$ (33)
with the optimal policy given by $\pi(a\mid s)=\exp\bigl((Q(s,a)-V(s))/\tau\bigr)$. The operator $\mathcal{T}^{\mathrm{soft}}$ defined by Equations (32)–(33) is a $\gamma$-contraction in $\|\cdot\|_{\infty}$, with the same convergence rate as the standard Bellman operator. [^46] The proof follows from the log-sum-exp being 1-Lipschitz in the sup norm; see Haarnoja et al. (2017), Theorem 1. As $\tau\to 0$, the log-sum-exp converges to the hard maximum and the soft Bellman equations reduce to the standard Bellman optimality equations. The value function in Equation (33) is identical to McFadden’s social surplus function from discrete choice theory (McFadden, 1978), completing the connection noted in the preceding subsection.
Classical RL already employs exploration strategies ($\varepsilon$-greedy, UCB), but these are designed separately from the optimization objective. In the probabilistic framework, the objective itself maximizes expected reward and policy entropy simultaneously. [^47] The max-ent objective is $\max_{\pi}\sum_{t}\mathbb{E}[r(s_{t},a_{t})+\tau\mathcal{H}(\pi(\cdot|s_{t}))]$. This equals the evidence lower bound (ELBO) of the graphical model, a quantity from variational inference that lower-bounds the log-likelihood of the observed optimality evidence $\log P(\mathcal{O}_{1:T}=1)$. Maximizing the ELBO with respect to the policy is equivalent to finding the best approximation to the true posterior, measured by KL divergence; see Blei et al. (2017) for a review of variational inference. The Q-function, the value function, and the exploration behavior are all derived from a single objective. Different approximation strategies applied to this single model recover familiar algorithms. When the backward messages in Equations (30)–(31) are computed exactly in the tabular setting, the result is soft value iteration. Estimating the ELBO gradient via the likelihood ratio trick yields max-ent policy gradients, identical to REINFORCE with $\tau\log\pi(a_{t}|s_{t})$ added to the reward at each step. Fitting parameterized $Q_{\phi}$ and $V_{\psi}$ networks to approximate the backward messages gives Soft Actor-Critic (Haarnoja et al., 2018). Fitting $Q_{\phi}$ alone and extracting the policy implicitly via the softmax gives soft Q-learning (Haarnoja et al., 2017). Trust-region methods such as TRPO and PPO do not optimize the max-ent objective, but each policy update step solves a max-ent subproblem with the old policy as prior, so the per-step structure mirrors the framework even though the global target remains standard reward maximization (Schulman et al., 2015, 2017). Hard-max algorithms (Q-learning, DQN, standard policy gradient) are not direct instances of the framework; they are recovered in the zero-temperature limit $\tau\to 0$, where the softmax collapses to the argmax. Reading the graphical model in the reverse direction, treating the reward as the unknown and observed behavior as evidence of optimality, yields maximum entropy inverse reinforcement learning (Ziebart et al., 2008), connecting to the structural estimation methods discussed in Section 8.
The framework has practical limitations. The converged fixed point is $Q^{*}_{\mathrm{soft}}$, not $Q^{*}$; at any $\tau>0$, the policy is suboptimal for the original reward-maximization objective. The entropy bonus produces undirected exploration, spreading probability mass rather than targeting uncertain states. In safety-critical environments, the objective assigns nonzero probability to catastrophic actions. The practical benefits of the probabilistic perspective, including robustness to model perturbations and smooth optimization landscapes, are most apparent in continuous-control and robotics settings. In perfectly simulated environments with exact rewards (Atari, Go), the hard-max algorithms that dominate those domains do not require this probabilistic formulation, as the following subsection illustrates.
@@ 5.2.8 Rollout, Lookahead, and AlphaZero
Recall the AlphaGo Zero system from Section 4.2.5, where a neural network $f_{\theta}(s)=(\mathbf{p},v)$ outputs a prior policy $\mathbf{p}$ and a value estimate $v$ for any board position $s$. During play, the network does not act alone: Monte Carlo Tree Search runs simulated games from the current position, using $v$ to evaluate leaf nodes and $\mathbf{p}$ to guide which branches to explore. [^81] AlphaGo Zero uses 1,600 MCTS simulations per move for Go; the generalized AlphaZero algorithm uses 800 simulations per move across chess, shogi, and Go (Silver et al., 2018a, Table S3). The network provides $\tilde{V}$; MCTS applies multi-step lookahead through selective tree expansion. Table 7 makes the correspondence explicit.
@@ 5.4.3 Natural Policy Gradient and Gradient Domination
The RL objective $J(\theta)$ is non-convex in $\theta$. For researchers trained to distrust gradient methods on non-convex objectives, the natural concern is convergence to spurious local optima. For tabular softmax policies (one free parameter per state-action pair), this concern is unfounded. The landscape is “benign” in a precise sense. Agarwal et al. (2021a) prove that $J(\theta)$ satisfies a gradient domination condition (also called Polyak-Łojasiewicz, or PL). The PL condition has the same functional form as the strong convexity condition for guaranteeing linear convergence of gradient descent, but it applies to non-convex functions: whenever $\|\nabla J(\theta)\|$ is small, the policy must be near-optimal. Formally, the sub-optimality $J(\pi^{*})-J(\pi_{\theta})$ is bounded by a constant times $\|\nabla J(\theta)\|^{2}$. The implication is immediate: any point where the gradient vanishes is globally optimal. The non-convex landscape has no false peaks, no spurious local maxima. Gradient ascent cannot get trapped.
@@ 5.5.1 Actor-Critic Architecture and Two-Timescale Convergence
Convergence requires a structural condition on the critic. The critic’s feature vectors must span the actor’s score functions $\nabla_{\omega}\log\pi_{\omega}(a|s)$, so that the critic’s approximation error lies orthogonal to the policy gradient direction. Under this compatibility condition, the critic’s projection error does not bias the actor’s gradient estimates.100100100The compatible function approximation theorem first appears in Sutton et al. (2000) and is the key structural requirement in Konda and Tsitsiklis (2000). It constrains the critic architecture to be “compatible” with the actor parameterization, the same condition that makes the natural policy gradient equal to the critic’s weight vector in Kakade (2002).
A2C (Advantage Actor-Critic) is the synchronous variant: collect a batch of transitions, compute TD errors, and update both networks. A3C (Mnih et al., 2016) parallelizes this across multiple workers updating a shared parameter server asynchronously.101101101Parallel workers decorrelate gradient estimates by exploring different parts of the state space simultaneously, removing the need for an experience replay buffer. Rigorous convergence theory for A3C’s lock-free asynchronous parameter updates remains an open problem; existing analyses of asynchronous stochastic approximation (Qu and Wierman, 2020) address classical asynchrony (different state-action pairs updated at different times on a single trajectory), not the parallel-worker gradient setting.
@@ 5.5.2 Entropy Regularization and Soft Actor-Critic
where $\mathcal{H}(\pi)=-\sum_{a}\pi(a)\log\pi(a)$ is the entropy and $\tau>0$ is the temperature parameter. Geist et al. (2019) later provided the unifying theoretical framework, showing that entropy regularization converts the Bellman optimality operator’s non-smooth hard max into a smooth log-sum-exp, and that the resulting soft Bellman operator remains a $\gamma$-contraction, preserving the convergence guarantees of standard dynamic programming.102102102The optimal policy under entropy regularization is $\pi^{*}(a|s)\propto\exp(Q^{*}(s,a)/\tau)$, which is precisely the McFadden (1974) logit choice probability with systematic utility $Q^{*}(s,a)$ and scale parameter $\tau$. The entropy-regularized value function is the log-sum-exp of Q-values, corresponding to the inclusive value (log-sum) operator in nested logit models. This equivalence between the soft-control framework and dynamic discrete choice models with EV1 taste shocks is developed formally in Rust and Rawat (2026), Appendix A.
@@ 5.5.3 Error Amplification Under Approximate Value Functions
At $\gamma=0.99$, the amplification factor is $2\cdot 0.99/0.01=198$. A 1% error in value function approximation yields at most 198% error in policy value.103103103The Singh-Yee bound is worst-case and not tight in general; massoud Farahmand et al. (2010) derive tighter bounds under smoothness assumptions on the MDP. The amplification factor $2\gamma/(1-\gamma)$ is a sensitivity analysis: it quantifies how errors in the “inputs” (value estimates) propagate to “outputs” (policy quality), analogous to errors-in-variables bias in regression. The discount factor $\gamma$ controls sensitivity; more patient agents face larger amplification. This bound is pessimistic but finite: approximate value functions do not cause unbounded policy degradation.
@@ 5.5.4 Sample Complexity of Planning
Classical dynamic programming complexity scales with the state space size $|\mathcal{S}|$. For problems like Go, where $|\mathcal{S}|\approx 10^{170}$, exact computation is impossible. Kearns et al. (2002) prove that with access to a generative model104104104A “generative model” in RL is a simulator that, given any state-action pair $(s,a)$, returns a sampled next state $s^{\prime}\sim P(\cdot|s,a)$ and reward $r$. This is unrelated to “generative models” in machine learning (GANs, diffusion models) or “generative processes” in Bayesian econometrics. The distinction matters: planning with a generative model is strictly easier than learning from a single trajectory, because the agent can query arbitrary states rather than following a sequential path. (a simulator that samples transitions from any state-action pair), near-optimal planning is possible with no dependence on $|\mathcal{S}|$. The cost is exponential dependence on the effective horizon $H=\log(R_{\max}/(\epsilon(1-\gamma)))/\log(1/\gamma)$: the sparse sampling algorithm requires $O((|\mathcal{A}|/\epsilon)^{H})$ simulator calls.105105105The sparse sampling algorithm builds a random tree of depth $H$ from the current state, sampling $C$ successor states per action at each node, then estimates values by averaging leaf rewards back up the tree. Its running time is $O((C|\mathcal{A}|)^{H})$, which is exponential in $H$ but entirely independent of $|\mathcal{S}|$. Kearns et al. (2002) also prove a lower bound of $\Omega(2^{H})$ generative model calls for any planning algorithm, so exponential horizon dependence is unavoidable in the worst case. For $\gamma$ near 1, $H\approx(1-\gamma)^{-1}\log(R_{\max}/\epsilon)$, so the method is practical only for short effective horizons or moderate discount factors. The key insight is the tradeoff: classical DP scales linearly in $|\mathcal{S}|$ but polynomially in $H$; sparse sampling eliminates state-space dependence at the cost of exponential horizon dependence. This explains why MCTS succeeds in large state spaces with bounded lookahead.
@@
The minimax-optimal sample complexity for planning with a generative model, when queries to arbitrary state-action pairs are permitted, is $\Theta(|\mathcal{S}||\mathcal{A}|/((1-\gamma)^{3}\epsilon^{2}))$ (Azar et al., 2013). This bound scales linearly in $|\mathcal{S}|$ but polynomially in $1/(1-\gamma)$, the opposite regime from sparse sampling. Agarwal et al. (2020a) show that the plug-in model-based approach (learn $\hat{P}$ from samples, then plan with $\hat{P}$) achieves this minimax rate, establishing that model-based RL is statistically optimal. Li et al. (2024b) further tighten this result by breaking the $|\mathcal{S}||\mathcal{A}|/(1-\gamma)^{2}$ sample-size barrier, showing that the minimax rate is achievable with total sample size as low as $|\mathcal{S}||\mathcal{A}|/(1-\gamma)$.106106106Recent extensions push these results beyond standard MDPs. Clavier et al. (2024) study the robust MDP setting where the agent must plan under model uncertainty with sa-rectangular or s-rectangular uncertainty sets, establishing minimax rates for robust policy optimization. Wang et al. (2025) extend the analysis to risk-sensitive objectives under the iterated CVaR criterion.